OpenStudy (anonymous):

3.04]What is the Least Common Denominator (LCD) of nine-sixteenths and nineteen twenty-fourths? forty-eight thirty-two sixteen 4 -------------------------------------------------------------------------------- Question 9 (Multiple Choice Worth 1 points) [3.05]When one U.S. Dollar is worth one point five two Australian Dollars, which of the following could be used as a conversion factor? nineteen one hundred twenty-fifths one and twelve twenty-fifths one and thirteen twenty-fifths five and one-fiftieth

OpenStudy (anonymous):

@Angel_1971

OpenStudy (anonymous):

need help

OpenStudy (stamp):

LCD\[\frac{9}{16},\ \frac{19}{24}\]\[16=2^4,\ 24=2^33^1\]\[2^4:2^33^1=16*3:24*2\]